OpenStudy (anonymous):

6(x+7)=4x-3 please show all the work :)

OpenStudy (anonymous):

6x+7 = 4x-3 6x - 4x= 2x (new equation equals= 4x+ 7= -3) -3 - 7= -10 4x=-10 -10/4= -2.5 x=-2.5